OpenStudy (anonymous):

prove: tan x/2 = the square root of 1-cosx/ 1+cosx

Parth (parthkohli):

\[\cos x = 2 \cos^2 \frac{x}2 - 1\]\[\cos x = 1 - 2\sin^2\frac{x}2\]Now find the values of \(1-\cos x \) and \(1 + \cos x\), then divide.

OpenStudy (welshfella):

- got beaten to it lol!

OpenStudy (anonymous):

thanks!!

Parth (parthkohli):

did you get it?

OpenStudy (anonymous):

yes i was just stuck on the second part and i have another problem that i am stuck on its prove: csc^2 x/2 = 2secx/secx -1

OpenStudy (anonymous):

@welshfella

OpenStudy (anonymous):

yes i have done that already

OpenStudy (welshfella):

sorry gtg right now

OpenStudy (anonymous):

okay thanks anyway

OpenStudy (welshfella):

im back

OpenStudy (welshfella):

heck gtg again

OpenStudy (welshfella):

convert to sines and cosines 1 / sin^2 (x/2) = (2/ cos x) / [(1/ cos x) - 1]

OpenStudy (welshfella):

now use Parthkholis second identity for cos x and plug it into the right side and simplify Its a bit messy but it will simplify to 1 / sin^2 (x/2)

OpenStudy (welshfella):